An old conveyor belt takes 21 h to move one day's coal output from the mine to a rail line. A new belt can do it in 15 h. How long does it take when both are used at the same time?
Please help. I've seriously spent about an hour trying to set this up and I can't, well I just haven't been able to yet.

Answer by richwmiller(17219)

x/21+x/15=1
x/3*7+x/3*5
5x/3*7*5+7x/3*5*7=1
5x/105+7x/105=1
12x/105=1
4x=35
x=35/4
x=8.75 hours